# MetaAligner: Conditional Weak-to-Strong Correction for Generalizable Multi-Objective Alignment of Language Models

### v1

@@ Abstract
Recent advancements in large language models (LLMs) aim to tackle heterogeneous human expectations and values via multi-objective preference alignment. However, existing methods are parameter-adherent to the policy model, leading to two key limitations: (1) the high-cost repetition of their alignment algorithms for each new target model; (2) they cannot expand to unseen objectives due to their static alignment objectives. In this work, we propose Meta-Objective Aligner (MetaAligner), a model that performs conditional weak-to-strong correction for weak responses to approach strong responses. MetaAligner is the first policy-agnostic and generalizable method for multi-objective preference alignment, which enables plug-and-play alignment by decoupling parameter updates from the policy models and facilitates zero-shot preference alignment for unseen objectives via in-context learning. Experimental results show that MetaAligner achieves significant and balanced improvements in multi-objective alignments on 11 policy models with up to 63$\times$ more parameters, and outperforms previous alignment methods with down to 22.27$\times$ less computational resources. The model also accurately aligns with unseen objectives, marking the first step towards generalizable multi-objective preference alignment. Models and codes will be available at: https://github.com/SteveKGYang/MetaAligner
@@ 1 Introduction
The recent advancement of large language models (LLMs) is largely reflected in their high-quality responses that align with human expectations and values. At the final stage of alignment, LLMs are supervised on human preference data via reinforcement learning from human feedback (RLHF) (Ziegler et al., 2019; Ouyang et al., 2022; Stiennon et al., 2020), where a proxy, directly trained on human preferences data, is leveraged to provide scalar rewards for reinforcement learning on the target model (Ouyang et al., 2022).
@@
RLHF (Ouyang et al., 2022) PPO
MORLHF (Li et al., 2020) PPO
MODPO (Zhou et al., 2023) DPO
CPO (Guo et al., 2024) SFT, DPO
RiC (Yang et al., 2024) SFT
Aligner (Ji et al., 2024) SFT
@@
However, human expectations and values include a broad spectrum of heterogeneous and multi-dimensional objectives, which makes scalar supervisions inefficient for aligning diverse and inclusive human preferences (Bai et al., 2022; Rame et al., 2024). These drawbacks motivate further exploration into multi-objective alignment algorithms. Some intuitive methods extend RLHF into multi-objective RLHF (MORLHF) (Sener and Koltun, 2018; Li et al., 2020; Rame et al., 2024). Due to its substantial computational cost (Li et al., 2020; Rame et al., 2024) and the unstable nature of the PPO (Schulman et al., 2017; Kumar et al., 2020; Rafailov et al., 2024), other methods seek to bypass the RL paradigm with multi-objective direct preference optimization (MODPO) (Zhou et al., 2023) or supervised fine-tuning (SFT)-based methods (Yang et al., 2024; Guo et al., 2024), which customized prompting strategies to incorporate multiple reward values into queries explicitly.
The above methods for multi-objective alignment bear one commonality: the parameter adherence to the policy model. This paradigm inevitably brings two key limitations: (1) they require repetition of their high-cost alignment algorithms for each newly-introduced target model, which is incompatible with the increasing sizes and fast iteration of current foundation models (Achiam et al., 2023; Team et al., 2023; Touvron et al., 2023; Chiang et al., 2023; Team et al., 2024). (2) all target models are statically aligned on pre-determined (e.g. Helpfulness, Harmlessness, Honesty (Zhou et al., 2023; Guo et al., 2024)) objectives, with currently no efforts in expanding and evaluating their capabilities on unseen objectives. This ignorance leads to poor generalizability of existing multi-objective alignment methods.
@@
In this work, we propose Meta-Objective Aligner (MetaAligner), the first policy-agnostic and generalizable method for multi-objective preference alignment. MetaAligner decouples preference alignment from the policy models and enables plug-and-play inferences, even without policy parameter access. This is achieved by extending the Aligner (Ji et al., 2024) to multi-objective alignment scenarios, which are not directly solvable by Aligner itself. Specifically, typical multi-objective alignment datasets are reorganized into dynamic multi-objective datasets, training the model to perform reward-free alignment and flexible adjustment of target objectives. The enhanced datasets supervise MetaAligner via conditional weak-to-strong correction. That is, MetaAligner is trained to correct the weak responses to approach the strong responses considering the corresponding alignment objectives, which are incorporated via prompt engineering. During inference, MetaAligner is stacked onto any policy models to perform objective-aware corrections. As MetaAligner auto-regressively attends to target objectives and requires no explicit reward values, it bears a high potential for expansion to unseen objectives, a new feature with rare previous exploration in LLM alignment. The model can be easily adapted to new objectives by updating the objective descriptions in the prompts, and new alignment strategies for correction are achieved via in-context learning (Kojima et al., 2022). The number of aligned objectives for MetaAligner also becomes expandable, theoretically leading to unlimited simultaneous alignment objectives. A checklist of MetaAligner’s key advantages over previous methods is presented in Table 1.
Through this work, we make three main contributions:
@@
  • We propose MetaAligner, the first policy-agnostic method for multi-objective preference alignment. It performs multi-objective alignment efficiently, without tuning the policy models or even accessing their parameters. Experimental results show that MetaAligner outperforms previous alignment methods with down to 22.27× less computational resources and higher stability.
  • We explore utilizing MetaAligner to exert zero-shot preference alignment for unseen objectives. To the best of our knowledge, this work marks the first attempt at generalizable multi-objective preference alignment. Experimental results show that MetaAligner performs effective alignment for 3 unseen objectives while maintaining performance on aligned objectives.
  • We examine MetaAligner on the IMHI benchmark (Yang et al., 2023b), which consists of 9 sub-tasks for mental health analysis and 20K testing samples. Experimental results show that MetaAligner trained on merely 7.2K human preference data improves win-rate on multiple objectives across 11 policy models with sizes ranging from 2B to 70B, substantially enhancing responses from policy models with up to 63$\times$ more parameters.
@@
Figure 1: Illustrations of three multi-objective preference alignment categories: (A) MORLHF (Li et al., 2020), (B) SFT-based methods (Yang et al., 2024; Guo et al., 2024), and (C) Meta-Objective Aligner. SFT-based methods are trained in a two-stage way and MetaAligner follows a three-stage paradigm.
2 Multi-Objective Alignment
@@
In real-world scenarios, human expectations of high-quality responses from AI agents involve considerable variability, with complex interplay such as contradiction (e.g. helpfulness and harmlessness) and dependence (e.g. correctness and informativeness). Multi-objective preference alignment tackles this challenge by optimizing on multiple objectives, which is formalized as: $\mathop{argmax}_{\pi_{\phi}}\mathbf{R}=[r_{1},...,r_{N}]^{\mathrm{T}}$, where $\pi_{\phi}$ denotes the aligned policy model parameterized by $\phi$. $r_{i}$ denotes the reward values for $i$-th objective, which is defined, in most cases of preference alignment, under the Bradley-Terry (Bradley and Terry, 1952) model of preferences. Specifically, for the same prompt $q$ and two responses $(y_{1},y_{2})$ under data distribution $\mathcal{D}$, the model assumes:
@@
where $P_{\mathcal{D}}(y_{1}\succ y_{2})$ denotes the probability that $y_{1}$ is preferred against $y_{2}$. MORLHF aims to achieve Pareto optimal among objectives, where the policy model is optimized to maximize a linear scalarization of multiple rewards (Sener and Koltun, 2018; Li et al., 2020) with a KL-divergence regularization (see Figure 1 (A)):
$\small\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{q\sim\mathcal{D},y\sim\pi_{\phi}}\left[\omega^{\mathrm{T}}\mathbf{R}(q,y)-\beta log\frac{\pi_{\phi}(y|q)}{\pi_{sft}(y|q)}\right]$ (2)
where $\omega=[\omega_{1},...,\omega_{N}]\ s.t.\sum_{i=1}^{N}\omega_{i}=1,\omega_{i}\geq 0$ is the heuristic target preference vector. In another paradigm, Yang et al. (2024); Guo et al. (2024) directly build alignment between multiple reward values and their corresponding response via SFT (see Figure 1 (B)):
@@ 3 Meta-Objective Aligner
This section introduces the proposed methodology for MetaAligner, which follows a three-stage paradigm: (1) Dynamic multi-objective dataset construction; (2) a three-step model training process; (3) policy-agnostic and generalizable inference for alignment. The paradigm is illustrated in Figure 1 (C).
@@ 3.1 Dynamic Multi-Objective Dataset
We propose constructing a dynamic multi-objective dataset to trigger MetaAligner’s ability for reward-free and flexible adjustment of alignment objectives. Specifically, a typical multi-objective preference alignment dataset with $N$ objectives $\mathcal{D}_{n}$ is formalized as $\{q_{i},y_{i1},y_{i2},R_{i}\}_{i=1}^{n}$, where $R_{i}=[r_{i1},...,r_{iN}]^{\mathrm{T}}$ and $r_{ij}\in\{\succ,\prec,=\}$. We define a text marker for each objective: $<R>:[<r_{1}>,...,<r_{N}>]$, where $<r_{j}>$ denotes a natural language description marking the $i$-th objective. Six examples of such text markers are presented in Figure 1 (C). With a pre-defined prompting template $\mathcal{T}(q,y,\mathcal{O},t)$, we build a preference subset $\mathcal{D}_{p}$ and another equal-preference subset $\mathcal{D}_{e}$ from $\mathcal{D}_{n}$, where $\mathcal{D}_{p}$ includes all contrastive response pairs and $\mathcal{D}_{e}$ includes all equal-labelled response pairs. We utilize the following prompting template for building the datasets:
@@
where $q$ denotes the query, $y$ denotes a corresponding response, $\mathcal{O}$ denotes the text markers for the target objectives, and $t\in\{equal,better\}$ depends on the building subset. The detailed construction process is described in Algorithm 1.
Algorithm 1 Training data construction
1:Raw datasets: $\{q_{i},y_{i1},y_{i2},R_{i}\}_{i=1}^{n}$; Objective text markers: $[<r_{1}>,...,<r_{N}>]$; Prompting template: $\mathcal{T}(q,y,\mathcal{O},t)$
2:Preference subset $\mathcal{D}_{p}$; All-equal subset $\mathcal{D}_{e}$.
3:$\mathcal{D}_{p}\leftarrow\varnothing$, $\mathcal{D}_{e}\leftarrow\varnothing$triangleright Initialize the 2 subsets.
4:for $i\in\{1,...,n\}$ do
5:    $\mathcal{O}_{\succ}\leftarrow\varnothing$, $\mathcal{O}_{\prec}\leftarrow\varnothing$, $\mathcal{O}_{=}\leftarrow\varnothing$
6:    for $j\in\{1,...,N\}$ do
7:         if $r_{ij}$ is $\succ$ $\triangleright$ Preferred obj. for $y_{1}$.
8:             $\mathcal{O}_{\succ}\leftarrow\mathcal{O}_{\succ}\cup\{<r_{j}>\}$
9:         else if $r_{ij}$ is $\prec$ $\triangleright$ Preferred obj. for $y_{2}$.
10:             $\mathcal{O}_{\prec}\leftarrow\mathcal{O}_{\prec}\cup\{<r_{j}>\}$
11:         $\triangleright$ Equally preferred objectives.
12:             $\mathcal{O}_{=}\leftarrow\mathcal{O}_{=}\cup\{<r_{j}>\}$
13:         end if
14:    end for
15:    if $\mathcal{O}_{\succ}\neq\varnothing$ $\triangleright$ Build $y_{1}$-preferred data.
16:         $t\leftarrow<better>$
17:         $\mathcal{O}_{\succ}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{\succ})$
18:         $\mathcal{D}_{p}\leftarrow\mathcal{D}_{p}\cup\{(\mathcal{T}(q_{i},y_{i2},\mathcal{O}_{\succ},t),y_{i1})\}$
19:    end if
20:    if $\mathcal{O}_{\prec}\neq\varnothing$ $\triangleright$ Build $y_{2}$-preferred data.
21:         $t\leftarrow<better>$
22:         $\mathcal{O}_{\prec}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{\prec})$
23:         $\mathcal{D}_{p}\leftarrow\mathcal{D}_{p}\cup\{(\mathcal{T}(q_{i},y_{i1},\mathcal{O}_{\prec},t),y_{i2})\}$
24:    end if
25:    if $\mathcal{O}_{=}\neq\varnothing$ $\triangleright$ Build equal data.
26:         $t\leftarrow<equal>$
27:         $\mathcal{O}_{=}\leftarrow\textit{random\_shuffle}(\mathcal{O}_{=})$
28:         $\mathcal{D}_{e}\leftarrow\mathcal{D}_{e}\cup\{(\mathcal{T}(q_{i},y_{i2},\mathcal{O}_{=},t),y_{i1})\}$
29:    end if
30:end for
Training on dynamic multi-objective datasets provides 3 key advantages: (1) instance-level alternation of the objectives during training enables flexible alignment on different combinations of conditions; (2) mutual alignment between the same response pairs on different objectives fully leverages the supervision information in the preference vectors; (3) the reward-free alignment method avoids complicated preference-to-reward mapping (Yang et al., 2024) process in previous SFT-based multi-objective alignment methods.
@@ 3.2 MetaAligner Derivation
The above action achieves policy-agnostic alignment, which poses three advantages: (1) any data points from the dynamic multi-objective dataset become eligible for supervision, where weak responses are directly regarded as output from unknown policy models; (2) the computation resources required for MetaAligner training is detached from policy model size, which enables cost-efficient alignment for large foundation models; (3) MetaAligner works only via outputs from the policy models, which allows free fine-tuning and inference for alignment on closed-source LLMs (Achiam et al., 2023; Team et al., 2023).
@@ 3.3 Model Training
In practice, we utilize an LLM as the base model for MetaAligner, which provides domain knowledge and strong reasoning ability to support the conditional weak-to-strong correction process. We propose a three-step paradigm based on the objective function in Eqn. 6:
@@
The "warm-up" stage trains the model in identical response pairs, a prelude proven effective in residual correction strategies (He et al., 2016; Ji et al., 2024) to familiarize the model with identity mapping and facilitate convergence. Specifically, we randomly sample a subset of the equal-preference subset $\mathcal{D}_{e}$ as training samples but set an identical target response for each data point: $y\equiv y_{0}$.
Step 2: equal-preference modelling.
Due to the contrastive nature of their learning paradigm, most previous preference alignment works focus on modeling the residuals between response pairs and ignore the equal-preference response pairs. However, in objectives such as Correctness and Honesty (Yang et al., 2023b; Li et al., 2023b), equal preferences are common and enclose useful information such as the principle components of preference modeling regarding each objective. Based on these intuitions, we introduce a novel equal-preference modeling stage to fine-tune the warmed-up MetaAligner on the equal-preference subset $\mathcal{D}_{e}$.
@@ 3.4 Generalizable Inference
During inference, MetaAligner achieves alignment following the sampling process as in Eqn. 4, where unaligned outputs, sampled from the target policy model, are used as the input for conditional weak-to-strong correction. With the prompting-based paradigm, the target objectives for MetaAligner also become expandable and generalizable, a key advantage over previous alignment methods (Zhou et al., 2023; Yang et al., 2024; Guo et al., 2024). The generalizability is two-fold: Firstly, users can manipulate the target objectives by adjusting combinations of text markers in the objective set $\mathcal{O}$. For example, in alignment with objectives 1 and 3, we can incorporate the corresponding text markers as follows:
$\mathcal{O}={\color[rgb]{0,0,1}r_{3}:<r_{3}>;r_{1}:<r_{1}>}$ (7)
Secondly, the prompt-based objectives statement enables flexible adjustment of text markers for existing objectives and injections of unseen objectives. Following the last example, we have two unseen alignment objectives $r_{1}^{*}:<r_{1}^{*}>$, $r_{2}^{*}:<r_{2}^{*}>$, and an updated text marker $<r_{1}^{*}>$ for aligned objective 1. We can perform zero-shot alignment on these objectives by adjusting $\mathcal{O}$ as follows:
$\mathcal{O}^{*}={\color[rgb]{0,0,1}r_{3}:<r_{3}>;r_{1}:}{\color[rgb]{1,0,0}<r_{1}^{*}>;r_{1}^{*}:<r_{1}^{*}>;r_{2}^{*}:<r_{2}^{*}>}$ (8)

### v2

# MetaAligner: Towards Generalizable Multi-Objective Alignment of Language Models

@@ Abstract
Recent advancements in large language models (LLMs) aim to tackle heterogeneous human expectations and values via multi-objective preference alignment. However, existing methods are parameter-adherent to the policy model, leading to two key limitations: (1) the high-cost repetition of their alignment algorithms for each new target model; (2) they cannot expand to unseen objectives due to their static alignment objectives. In this work, we propose Meta-Objective Aligner (MetaAligner), a model that performs conditional weak-to-strong correction for weak responses to approach strong responses. MetaAligner is the first policy-agnostic and generalizable method for multi-objective preference alignment, which enables plug-and-play alignment by decoupling parameter updates from the policy models and facilitates zero-shot preference alignment for unseen objectives via in-context learning. Experimental results show that MetaAligner achieves significant and balanced improvements in multi-objective alignments on 10 state-of-the-art policy models, and outperforms previous alignment methods with down to 15.71$\times$ less GPU training hours. The model also effectively aligns unseen objectives, marking the first step towards generalizable multi-objective preference alignment. This project is open-sourced at: https://github.com/SteveKGYang/MetaAligner
@@ 1 Introduction
The recent advancement of large language models (LLMs) is largely reflected in their high-quality responses that align with human expectations and values. At the final stage of alignment, LLMs are supervised on human preference data via reinforcement learning from human feedback (RLHF) [39, 23, 28], where a proxy, directly trained on human preferences data, is leveraged to provide scalar rewards for reinforcement learning on the target model [23].
However, human expectations and values include a broad spectrum of heterogeneous and multi-dimensional objectives, which makes scalar supervisions inefficient for aligning diverse and inclusive human preferences [4, 25]. These drawbacks motivate further exploration into multi-objective alignment algorithms. Some intuitive methods extend RLHF into multi-objective RLHF (MORLHF) [27, 20, 25]. Due to its substantial computational cost [20, 25] and the unstable nature of the PPO [26, 17, 24], other methods seek to bypass the RL paradigm with multi-objective direct preference optimization (MODPO) [38, 11] or supervised fine-tuning (SFT)-based methods [35, 11], which customized prompting strategies to incorporate multiple reward values into queries explicitly.
The above methods for multi-objective alignment bear one commonality: the parameter adherence to the policy model. This paradigm inevitably brings two key limitations: (1) they require repetition of their high-cost alignment algorithms for each newly-introduced target model, which is incompatible with the increasing sizes and fast iteration of current foundation models [2, 31, 7, 30]; (2) all target models are statically aligned on pre-determined (e.g. "Helpful", "Harmless", "Honest" [38, 11]) objectives, with currently no efforts in expanding and evaluating their capabilities on unseen objectives. This ignorance leads to poor generalizability of existing multi-objective alignment methods.
@@
RLHF [23] PPO
MORLHF [20] PPO
MODPO [11, 38] SFT, DPO
RiC [35] SFT
Aligner [13] SFT
@@
In this work, we propose Meta-Objective Aligner (MetaAligner), the first policy-agnostic and generalizable method for multi-objective preference alignment. MetaAligner decouples preference alignment from the policy models and enables plug-and-play inferences, even without policy parameter access [13]. Specifically, typical multi-objective alignment datasets are reorganized into dynamic multi-objective datasets, training the model to perform reward-free alignment and flexible adjustment of target objectives. The enhanced datasets supervise MetaAligner via conditional weak-to-strong correction. That is, MetaAligner is trained to correct the weak responses to approach the strong responses considering the corresponding alignment objectives, which are incorporated via prompt engineering. During inference, MetaAligner is stacked onto any policy models to perform objective-aware corrections. As MetaAligner auto-regressively attends to target objectives and requires no explicit reward values, it bears a high potential for expansion to unseen objectives, a new feature with rare previous exploration in LLM alignment. The model can be easily adapted to new objectives by updating the objective descriptions in the prompts, and new alignment strategies for correction are achieved via in-context learning [16]. The number of aligned objectives for MetaAligner also becomes expandable, theoretically leading to unlimited simultaneous alignment objectives. A checklist of MetaAligner’s advantages over previous methods is presented in Table 1.
@@
Through this work, we make three main contributions: (1) we propose MetaAligner, the first policy-agnostic method for multi-objective preference alignment. It performs multi-objective alignment efficiently, without tuning the policy models or accessing their parameters. Experimental results show that MetaAligner outperforms previous alignment methods with down to 15.71× less GPU training hours and higher stability; (2) we utilize MetaAligner to exert zero-shot preference alignment for unseen objectives. To our knowledge, this work marks the first attempt at generalizable multi-objective preference alignment. Experimental results show that MetaAligner can simultaneously perform effective alignment for 6 unseen objectives while maintaining performance on aligned objectives; (3) We examine MetaAligner on 3 preference alignment datasets. Experimental results show that MetaAligner improves win rates on multiple objectives across 10 policy models, substantially enhancing responses of state-of-the-art foundation models such as GPT-3.5 [22] and Claude-3 [3].
@@
Figure 1: Illustrations of three multi-objective preference alignment categories: (A) MORLHF [20], (B) SFT-based methods [35, 11], and (C) Meta-Objective Aligner. SFT-based methods are trained in a two-stage way and MetaAligner follows a three-stage paradigm.
2 Multi-Objective Alignment of Language Models
@@
In real-world scenarios, human expectations of high-quality responses from AI agents involve considerable variability, with complex interplays such as contradiction (e.g. "Helpful" and "Harmless") and dependence (e.g. "Correct" and "Informative"). Multi-objective preference alignment tackles this challenge by optimizing multiple objectives, which is formalized as: $\mathop{argmax}_{\pi_{\phi}}\mathbf{R}=[r_{1},...,r_{N}]^{\mathrm{T}}$, where $\pi_{\phi}$ denotes the aligned policy model parameterized by $\phi$. $r_{i}$ denotes the reward values for $i$-th objective, which is defined, in most cases of preference alignment, under the Bradley-Terry [6] model of preferences. Specifically, for the same prompt $q$ and two responses $(y_{1},y_{2})$ under data distribution $\mathcal{D}$, the model assumes:
@@
where $P_{\mathcal{D}}(y_{1}\succ y_{2})$ denotes the probability that $y_{1}$ is preferred against $y_{2}$. MORLHF aims to achieve Pareto optimal among objectives, where the policy model is optimized to maximize a linear scalarization of multiple rewards [27, 20] with a KL-divergence regularization (see Figure 1 (A)):
$\mathop{argmax}\limits_{\pi_{\phi}}\mathbb{E}_{q\sim\mathcal{D},y\sim\pi_{\phi}}\left[\omega^{\mathrm{T}}\mathbf{R}(q,y)-\beta log\frac{\pi_{\phi}(y|q)}{\pi_{sft}(y|q)}\right]$ (2)
where $\omega=[\omega_{1},...,\omega_{N}]\ s.t.\sum_{i=1}^{N}\omega_{i}=1,\omega_{i}\geq 0$ is the heuristic target preference vector. Another paradigm [35, 11] directly built alignment between multiple reward values and their corresponding response via SFT (see Figure 1 (B)):
@@ 3 Meta-Objective Aligner
This section introduces the proposed methodology for MetaAligner, which follows a 3-stage paradigm: (1) dynamic multi-objective dataset construction; (2) a 3-step model training process; (3) generalizable inference for alignment. The paradigm is illustrated in Figure 1 (C).
@@ 3.1 Dynamic Multi-Objective Dataset
[…]
@@
where $q$ denotes the query, $y$ denotes a corresponding response, $\mathcal{O}$ denotes the text descriptions for the target objectives, and $t\in\{equal,better\}$ depends on the building subset. The detailed construction process is described in Algorithm 1. A full list of the utilized prompting templates is presented in Appendix B.
Training on dynamic multi-objective datasets provides 3 key advantages: (1) instance-level alternation of the objectives during training enables flexible alignment on different combinations of conditions; (2) mutual alignment between the same response pairs on different objectives fully leverages the supervision information in the preference vectors; (3) the reward-free alignment method avoids complicated preference-to-reward mapping [35] process in previous SFT-based multi-objective alignment methods.
@@ 3.2 MetaAligner Derivation
The above action achieves policy-agnostic alignment, which poses three advantages: (1) any data points from the dynamic multi-objective dataset become eligible for supervision, where weak responses are directly regarded as samples from unknown policy models; (2) the computation resources required for MetaAligner training is detached from policy model size, which enables cost-efficient alignment for large foundation models; (3) MetaAligner works only via outputs from the policy models, which allows free fine-tuning and inference for alignment on closed-source LLMs [2, 22, 3].
@@ 3.3 Model Training
In practice, we utilize an LLM as the base model for MetaAligner, which provides domain knowledge and strong reasoning ability to support the conditional weak-to-strong correction process. We propose a 3-step paradigm based on the objective function in Eqn. 6:
@@
The "warm-up" stage trains the model in identical response pairs, a prelude proven effective in residual correction strategies [12, 13] to familiarize the model with identity mapping and facilitate convergence. Specifically, we randomly sample a subset of the equal-preference subset $\mathcal{D}_{e}$ as training samples but set an identical target response for each data point: $y\equiv y_{0}$.
Step 2: equal-preference modeling.
Due to the contrastive nature of their learning paradigm, most previous preference alignment works focus on modeling the residuals between response pairs and ignore the equal-preference response pairs. However, equal preferences are common in many scenarios [34, 8] and enclose useful information such as the principle components of preference modeling regarding each objective. Based on these intuitions, we introduce a novel equal-preference modeling stage to fine-tune the warmed-up MetaAligner on the equal-preference subset $\mathcal{D}_{e}$.
@@ 3.4 Generalizable Inference
[…]